# Exposía: Academic Writing Assessment of Exposés and Peer Feedback

### v1

@@ G.2 Human–human agreement
$\alpha$ values range between -1 and 1, with 1 being perfect agreement, 0 being agreement by chance, and negative values indicating systematic disagreement; $r$ measures the correlation between a system’s predicted scores and the annotator-assigned scores; it ranges from -1 to 1. A positive (negative) value implies that the two sets of predictions are positively (negatively) correlated. Inter-annotator agreement results for exposé criteria are reported in Table 19, while agreement for review criteria is summarized in Tables 20 and 21.
@@ G.3 Failure modes of chance-corrected agreement
A closely related prevalence-driven failure mode appears in review scoring for error and penalty criteria (Tables 20 and 21). These criteria are effectively rare-event detection tasks on a highly skewed label space: the non-penalizing default (typically 0) dominates, while penalizing assignments (e.g., 1-1 in Errors (ER)) occur only rarely. As a result, chance-corrected reliability and Pearson correlation become unstable or uninformative. A small number of non-overlapping penalizations can yield $\alpha\leq 0$ despite high raw agreement, and near-constant rating vectors drive $r$ toward zero or render it undefined when one rater shows zero variance.
@@
Table 22: Quadratic-weighted agreement between LLM scores and human ratings for exposé criteria. Columns report $A_{G1}$ (Group 1 vs. LLM) and $A_{G2}$ (Group 2 vs. LLM) for the prompting variants shown (all, ind). Bold/underlined cells mark column-wise maxima/minima. Average rows summarize agreement within rubrics and across all criteria.
@@
Table 23: Quadratic-weighted agreement between LLM scores and human ratings for review criteria. Columns report $A_{G1}$ (Group 1 vs. LLM) and $A_{G2}$ (Group 2 vs. LLM) for the prompting variants shown (all, ind). Bold/underlined cells mark column-wise maxima/minima. Average rows summarize agreement within rubrics and across all criteria.
Appendix H Experimental Details
@@
We evaluate a diverse set of recent open-weight instruction-tuned LLMs that cover different architectures and scales. Our selection includes (i) a large dense decoder-only Transformer from the Llama 3.3 family (meta-llama/Llama-3.3-70B-Instruct; 70B parameters) Team (2024), (ii) a high-sparsity Mixture-of-Experts (MoE) model with long-context support (Qwen/Qwen3-Next-80B-A3B-Instruct; 80B total parameters with 3B active) Team (2025) and an independent MoE reasoning model in two flavors, (iii) the high reasoning model (openai/gpt-oss-120b, 117B parameters with 5.1B active parameters) and (iv) the smaller low latency variant (openai/gpt-oss-20b, 21B parameters with 3.6B active parameters) OpenAI (2025).
All models are fully open-weight and can be run on two local NVIDIA A100 80GB GPUs, which is important for classroom deployment where data protection and limited consent often require on-premise inference instead of cloud-based APIs. Inference is performed using the Hugging Face Transformers library [^15] https://huggingface.co/docs/transformers with deterministic greedy decoding (do_sample = False).
@@
For both tasks, we use a unified prompting strategy (Appendix I) with a system message that (i) assigns the model the role of an expert evaluator and (ii) enumerates the relevant criteria, including their names, short descriptions, allowed integer score ranges, and the verbal descriptions associated with each score level (Tables LABEL:tab:expose_rubric and LABEL:tab:review_rubric). The model is required to return a single JSON object with fields criterion_name, assigned_score, and a short justification. The additional justifications are requested to explain the decision-making process, as it is crucial in an environment where people are affected by the LLM outcome Kolářová and Schmude (2026).
@@
Appendix I Prompt Templates
@@
I.1 Expose Scoring Prompt
@@
I.2 Review Scoring Prompt
@@
Appendix J Additional Experimental Results
Agreement by expertise level
Table 4 reports human–LLM agreement (QWA) for criterion-level scoring, grouped by criterion expertise. Across both tasks, agreement decreases monotonically with expertise, mirroring the human–human reference. For exposé scoring, low-expertise criteria (Exp. 0; three criteria) yield high agreement for the strongest models under combined prompting (e.g., Llama 3.3: $\overline{A}_{\text{all}}{=}0.90$; Qwen3: $0.91$), whereas Exp. 2 criteria are substantially harder (Llama 3.3: $0.63$; Qwen3: $0.57$); humans show the same drop in the Human column (0.89 at Exp. 0 vs. 0.64 at Exp. 2), consistent with higher-expertise criteria being more subjective. Aggregated over all 31 exposé criteria, model agreement is close to the human baseline (Human: 0.76), and combined prompting is typically comparable or slightly better than single-criterion prompting, with Qwen3 as the main exception ($\overline{A}_{\text{ind}}{=}0.75$ vs. $\overline{A}_{\text{all}}{=}0.72$). The comparatively lower Exp. 0 values for GPT-OSS in the aggregate are explained by the small size of this bin and a single outlier criterion: Table 22 shows that GPT-OSS matches humans very well on the schedule checks (SC1/SC3; QWA $\approx 0.91$$0.98$), but performs unusually poorly on BI1 (Bibliography Consistency; e.g., GPT-OSS 120B: $A^{\text{all}}_{G1}{=}0.27$, $A^{\text{all}}_{G2}{=}0.40$), which dominates the Exp. 0 average. For review scoring, overall agreement is higher (Human: 0.88 over 22 criteria), and Llama 3.3 and Qwen3 reach near-human alignment under combined prompting (Llama 3.3: $\overline{A}_{\text{all}}{=}0.90$; Qwen3: $0.89$); the expertise effect persists, with Exp. 2 criteria yielding lower agreement than Exp. 1.
Figure 7: Human–LLM agreement (QWA) for review scoring by expertise level. The dotted line shows human-human QWA.
Figure 8: Group-asymmetry in human–LLM agreement by review rubric. Each point shows the difference in agreement between an LLM and human raters (Group 1 (G1) vs. Group 2 (G2)). The vertical line at $\Delta=0$ indicates equal agreement of the LLM with human G1 and G2 raters. Points to the right indicate the model agrees more with G1 raters than with G2 raters.
Figure 9: Runtime per submission for criterion-based assessments across models. Bars show mean seconds per submission with standard deviation bars. Within each task, All denotes combined prompting (light fill) and Ind denotes single-criterion prompting (dark fill). Review runs are hatched; Exposé runs are unhatched. Ratios show $\times(\mathrm{Ind}/\mathrm{All})$, highlighting the additional latency introduced by single-criterion prompting.
@@
Across all models and both tasks, combined prompting is more efficient in most cases than single-criterion prompting: it reduces mean runtime per submission by about $1.3\times$$5.3\times$ and avoids the linear cost growth from issuing one model call per criterion (Figure 9). Single-criterion prompting also substantially increases total token processing because the full input is repeated across prompts, with particularly large overhead for long exposés and the review setting. These efficiency differences become especially consequential under the observed course usage dynamics. As shown by the behavioural interaction logs (Appendix Figure 6), student activity is strongly deadline-driven, with a pronounced spike in actions immediately before the review deadline, whereas instructor activity is relatively steady. This implies that in practice, LLM-based scoring must be provisioned for short periods of peak demand rather than average-day usage. As such, the overhead of single-criterion prompting is particularly problematic, while combined prompting better supports batching and throughput during deadline-induced bursts. Consequently, combined prompting is the recommended operating point for classroom deployment under constrained on-premise resources, with model choice primarily determining the remaining latency–quality trade-off. Runtime and token usage are reported in Appendix K (Tables 2425).

### v2

@@ G.1 Raters, training, and scoring design
We refer to instructors who graded during the semester as Group 1. They attended weekly sessions and discussed edge cases. Group 1 raters were instructed to resolve uncertainty in favor of the student, because grades were at stake. Group 2 includes a different subset of instructors (with one who also participated in Group 1). They received the same guidelines, but they were not instructed to resolve uncertainty in favor of the student.
@@
Exposé (#31) Review (#22)
Rater / Model All Ind. All Ind.
Human raters 0.76 0.88
Claude Opus 4.6 0.80 0.79 0.90 0.85
Gemini 3.1 Pro 0.78 0.79 0.90 0.89
GPT-5.2 0.80 0.78 0.87 0.79
GPT OSS 120B 0.77 0.75 0.78 0.72
GPT OSS 20B 0.77 0.76 0.74 0.69
Llama 3.3 70B 0.76 0.75 0.90 0.82
Qwen3 80B 0.72 0.75 0.89 0.84
Table 17: Quadratic weighted agreement (QWA) for the two tasks, reported over all criteria only. Prompting variants: all (combined) and ind. (individual). Each LLM value shows human–LLM QWA averaged over both human raters and all criteria. The number of criteria are shown in the title (for exposés and reviews respectively). Bold indicates the best model performance per column.
@@ G.2 Human–human agreement
$\alpha$ values range between -1 and 1, with 1 being perfect agreement, 0 being agreement by chance, and negative values indicating systematic disagreement; $r$ measures the correlation between a system’s predicted scores and the annotator-assigned scores; it ranges from -1 to 1. A positive (negative) value implies that the two sets of predictions are positively (negatively) correlated. Inter-annotator agreement results for exposé criteria are reported in Table 18, while agreement for review criteria is summarized in Tables 19 and 20.
@@ G.3 Failure modes of chance-corrected agreement
A closely related prevalence-driven failure mode appears in review scoring for error and penalty criteria (Tables 19 and 20). These criteria are effectively rare-event detection tasks on a highly skewed label space: the non-penalizing default (typically 0) dominates, while penalizing assignments (e.g., 1-1 in Errors (ER)) occur only rarely. As a result, chance-corrected reliability and Pearson correlation become unstable or uninformative. A small number of non-overlapping penalizations can yield $\alpha\leq 0$ despite high raw agreement, and near-constant rating vectors drive $r$ toward zero or render it undefined when one rater shows zero variance.
@@
Table 21: Quadratic-weighted agreement between open-weight LLM scores and human ratings for exposé criteria. Columns report $A_{G1}$ (Group 1 vs. LLM) and $A_{G2}$ (Group 2 vs. LLM) for the prompting variants shown (all, ind). Bold/underlined cells mark column-wise maxima/minima. Average rows summarize agreement within rubrics and across all criteria.
Claude Opus 4-6 Gemini 3.1 Pro GPT-5.2
Code $A^{\text{all}}_{\text{G1}}$ $A^{\text{ind}}_{\text{G1}}$ $A^{\text{all}}_{\text{G2}}$ $A^{\text{ind}}_{\text{G2}}$ $A^{\text{all}}_{\text{G1}}$ $A^{\text{ind}}_{\text{G1}}$ $A^{\text{all}}_{\text{G2}}$ $A^{\text{ind}}_{\text{G2}}$ $A^{\text{all}}_{\text{G1}}$ $A^{\text{ind}}_{\text{G1}}$ $A^{\text{all}}_{\text{G2}}$ $A^{\text{ind}}_{\text{G2}}$
LA1 0.84 0.83 0.87 0.86 0.85 0.85 0.80 0.85 0.86 0.83 0.91 0.88
LA2 0.92 0.92 0.87 0.86 0.89 0.85 0.76 0.76 0.90 0.91 0.80 0.86
Language 0.88 0.87 0.87 0.86 0.87 0.85 0.78 0.81 0.88 0.87 0.85 0.87
ME1 0.93 0.92 0.88 0.89 0.92 0.94 0.84 0.83 0.90 0.89 0.89 0.88
Metadata 0.93 0.92 0.88 0.89 0.92 0.94 0.84 0.83 0.90 0.89 0.89 0.88
MO1 0.87 0.87 0.82 0.82 0.87 0.84 0.82 0.81 0.87 0.80 0.82 0.71
MO2 0.89 0.91 0.78 0.81 0.88 0.87 0.77 0.77 0.89 0.87 0.84 0.75
MO3 0.98 0.98 0.93 0.93 0.98 0.98 0.93 0.90 0.98 0.98 0.93 0.93
MO4 0.82 0.82 0.84 0.84 0.82 0.81 0.84 0.81 0.82 0.82 0.84 0.84
MO5 0.76 0.76 0.58 0.58 0.76 0.71 0.58 0.54 0.75 0.75 0.56 0.56
MO6 0.93 0.92 0.90 0.91 0.88 0.92 0.82 0.84 0.91 0.90 0.89 0.90
MO7 0.60 0.56 0.69 0.65 0.71 0.55 0.58 0.67 0.60 0.45 0.69 0.65
Motivation 0.84 0.83 0.79 0.79 0.84 0.81 0.76 0.76 0.83 0.80 0.79 0.76
AP1 0.91 0.93 0.87 0.93 0.91 0.95 0.87 0.83 0.91 0.89 0.87 0.85
AP2 0.91 0.90 0.85 0.89 0.88 0.87 0.79 0.74 0.92 0.89 0.89 0.78
AP3 0.80 0.84 0.77 0.85 0.75 0.80 0.67 0.70 0.88 0.85 0.84 0.79
AP4 0.84 0.82 0.67 0.69 0.80 0.78 0.64 0.66 0.80 0.84 0.67 0.67
AP5 0.76 0.62 0.67 0.67 0.78 0.68 0.65 0.66 0.80 0.73 0.64 0.71
AP6 0.85 0.85 0.85 0.82 0.80 0.79 0.77 0.81 0.87 0.88 0.87 0.84
AP7 0.67 0.67 0.67 0.64 0.69 0.66 0.69 0.80 0.73 0.71 0.69 0.64
AP8 0.96 0.96 0.85 0.85 0.96 0.98 0.85 0.84 0.96 0.96 0.85 0.85
Approach 0.84 0.82 0.78 0.79 0.82 0.81 0.74 0.75 0.86 0.84 0.79 0.77
MD1 0.73 0.76 0.64 0.64 0.71 0.68 0.55 0.56 0.65 0.69 0.64 0.64
MD2 0.78 0.78 0.73 0.69 0.76 0.71 0.71 0.71 0.78 0.75 0.73 0.69
MD3 0.85 0.73 0.65 0.64 0.84 0.79 0.71 0.67 0.85 0.80 0.69 0.71
MD4 0.65 0.62 0.71 0.75 0.62 0.54 0.75 0.83 0.69 0.58 0.71 0.75
MD5 0.84 0.80 0.69 0.73 0.85 0.66 0.67 0.59 0.71 0.65 0.67 0.76
MD6 0.69 0.71 0.76 0.75 0.71 0.68 0.60 0.66 0.71 0.73 0.75 0.55
MD7 0.65 0.62 0.47 0.44 0.62 0.71 0.44 0.49 0.64 0.64 0.67 0.49
Methodology 0.74 0.72 0.66 0.66 0.73 0.68 0.63 0.64 0.72 0.69 0.69 0.65
SC1 0.96 0.98 0.96 0.98 0.96 0.97 0.96 0.93 0.96 0.96 0.96 0.96
SC2 0.98 1.00 0.95 0.96 0.98 0.97 0.95 0.93 0.98 0.98 0.95 0.95
SC3 0.91 0.91 0.96 0.96 0.91 0.93 0.96 1.00 0.91 0.93 0.96 0.98
SC4 0.84 0.90 0.91 0.89 0.80 0.85 0.78 0.85 0.89 0.89 0.95 0.91
Schedule 0.92 0.95 0.95 0.95 0.91 0.93 0.91 0.92 0.94 0.94 0.95 0.95
BI1 0.58 0.25 0.56 0.35 0.62 0.62 0.67 0.62 0.27 0.15 0.40 0.35
BI2 0.87 0.83 0.86 0.90 0.92 0.90 0.72 0.72 0.86 0.89 0.87 0.84
Bibliography 0.72 0.54 0.71 0.62 0.77 0.76 0.70 0.67 0.57 0.52 0.63 0.59
All criteria 0.83 0.81 0.78 0.78 0.82 0.80 0.75 0.75 0.82 0.79 0.79 0.76
Table 22: Quadratic-weighted agreement between closed-source LLM scores and human ratings for exposé criteria. Columns report $A_{G1}$ (Group 1 vs. LLM) and $A_{G2}$ (Group 2 vs. LLM) for the prompting variants shown (all, ind). Bold/underlined cells mark column-wise maxima/minima. Average rows summarize agreement within rubrics and across all criteria.
@@
Table 23: Quadratic-weighted agreement between open-weight LLM scores and human ratings for review criteria. Columns report $A_{G1}$ (Group 1 vs. LLM) and $A_{G2}$ (Group 2 vs. LLM) for the prompting variants shown (all, ind). Bold/underlined cells mark column-wise maxima/minima. Average rows summarize agreement within rubrics and across all criteria.
Claude Opus 4-6 Gemini 3.1 Pro GPT-5.2
Code $A^{\text{all}}_{\text{G1}}$ $A^{\text{ind}}_{\text{G1}}$ $A^{\text{all}}_{\text{G2}}$ $A^{\text{ind}}_{\text{G2}}$ $A^{\text{all}}_{\text{G1}}$ $A^{\text{ind}}_{\text{G1}}$ $A^{\text{all}}_{\text{G2}}$ $A^{\text{ind}}_{\text{G2}}$ $A^{\text{all}}_{\text{G1}}$ $A^{\text{ind}}_{\text{G1}}$ $A^{\text{all}}_{\text{G2}}$ $A^{\text{ind}}_{\text{G2}}$
SC1 0.85 0.84 0.89 0.87 0.86 0.87 0.86 0.87 0.84 0.85 0.86 0.86
SC2 0.90 0.85 0.88 0.88 0.90 0.90 0.84 0.89 0.83 0.87 0.85 0.85
Structure and Clarity 0.87 0.85 0.88 0.87 0.88 0.88 0.85 0.88 0.84 0.86 0.86 0.85
LG1 0.81 0.80 0.80 0.83 0.85 0.81 0.93 0.91 0.91 0.83 0.92 0.89
LG2 0.98 0.99 0.99 0.99 0.98 0.99 0.99 0.99 0.98 0.98 0.99 0.98
Language 0.90 0.90 0.90 0.91 0.92 0.90 0.96 0.95 0.94 0.91 0.95 0.94
FU1 0.87 0.87 0.87 0.83 0.78 0.83 0.76 0.77 0.83 0.85 0.80 0.79
FU2 0.87 0.84 0.85 0.82 0.78 0.80 0.75 0.76 0.83 0.84 0.79 0.79
Feed Up 0.87 0.86 0.86 0.83 0.78 0.81 0.76 0.77 0.83 0.84 0.79 0.79
FB1 0.94 0.75 0.99 0.75 0.94 0.75 0.99 0.76 0.94 0.65 0.99 0.64
FB2 0.93 0.75 0.94 0.74 0.94 0.90 0.95 0.96 0.93 0.28 0.92 0.22
FB3 0.78 0.75 0.74 0.71 0.74 0.60 0.71 0.62 0.62 0.42 0.65 0.47
FB4 0.84 0.85 0.96 0.97 0.84 0.83 0.96 0.99 0.84 0.84 0.96 0.96
FB5 0.95 0.95 0.94 0.94 0.95 0.96 0.94 0.94 0.94 0.95 0.93 0.94
FB6 0.95 0.97 0.96 0.98 0.99 0.99 0.96 0.97 0.97 0.85 0.94 0.84
Feed Back 0.90 0.84 0.92 0.85 0.90 0.84 0.92 0.87 0.88 0.66 0.90 0.68
FF1 0.96 0.96 0.99 0.99 0.96 0.97 0.99 0.99 0.95 0.96 0.98 0.99
FF2 0.81 0.79 0.68 0.61 0.81 0.73 0.70 0.69 0.79 0.79 0.68 0.63
Feed Forward 0.89 0.88 0.83 0.80 0.89 0.85 0.84 0.84 0.87 0.88 0.83 0.81
CQ1 0.92 0.86 0.92 0.89 0.96 0.96 0.92 0.92 0.80 0.80 0.83 0.83
CQ2 0.92 0.94 0.94 0.93 0.94 0.93 0.92 0.91 0.94 0.95 0.92 0.92
CQ3 0.97 0.98 0.99 0.99 0.98 0.98 0.99 0.99 0.98 0.98 0.99 0.99
Content Quality 0.94 0.92 0.95 0.93 0.96 0.96 0.94 0.94 0.91 0.91 0.92 0.91
ER1 0.86 0.84 0.91 0.85 0.84 0.84 0.89 0.88 0.75 0.79 0.79 0.83
ER2 0.85 0.84 0.85 0.82 0.93 0.97 0.93 0.92 0.85 0.90 0.83 0.88
ER3 0.94 0.84 0.92 0.79 0.95 0.85 0.96 0.84 0.86 0.81 0.85 0.78
ER4 0.85 0.42 0.87 0.38 0.94 0.93 0.98 0.98 0.79 0.22 0.81 0.18
ER5 0.99 0.99 0.99 0.99 0.99 0.91 0.99 0.91 0.99 0.96 0.99 0.96
Errors 0.90 0.78 0.91 0.77 0.93 0.90 0.95 0.91 0.85 0.73 0.85 0.73
All criteria 0.90 0.85 0.90 0.84 0.90 0.88 0.91 0.88 0.87 0.79 0.88 0.78
Table 24: Quadratic-weighted agreement between closed-source LLM scores and human ratings for review criteria. Columns report $A_{G1}$ (Group 1 vs. LLM) and $A_{G2}$ (Group 2 vs. LLM) for the prompting variants shown (all, ind). Bold/underlined cells mark column-wise maxima/minima. Average rows summarize agreement within rubrics and across all criteria.
Appendix H Illustrative Examples
This chapter presents two representative examples from Exposía. The examples illustrate how exposé drafts, reviewer comments, and revised final versions are connected in the dataset. Each example is shown in three parts: an annotated draft excerpt, a short summary of the reviewer feedback, and the corresponding final excerpt. Passages highlighted in red indicate text spans that received inline comments from the reviewer, in this case, marking weaknesses that were addressed during revision.
Example 1: Anchor
This example illustrates how reviewers ask students to strengthen the anchor of a proposal by making the motivation more concrete and evidence-based. In the draft, the problem context is understandable, but still too general: the text claims that museums have experienced a major decline in visitor numbers and need more engaging approaches, yet the argument is not substantiated with references or concrete examples. This is reflected not only in the human assessment, but also in the automatic scoring of the draft, where models differed in whether they interpreted the problem description as already sufficiently detailed or still somewhat too general. The automatic assessments of the draft show a clear split, which helps explain why the human evaluation was not entirely uniform either. First, most models already viewed the problem statement as sufficiently detailed. For example, Claude Opus 4.6 assigned a score of 2 and argued that “the problem is described in detail,” explicitly pointing to the post-COVID decline in visitor numbers, the reliance on traditional text-panel-based interaction, and the limited current use of holograms. Second, a smaller group of models judged the same passage more critically because the motivation remained somewhat general. For instance, GPT OSS (20B) assigned a score of 1 and summarized the draft as “Problem stated with some detail but missing deeper specifics.”
The student review is especially informative when viewed through Feed Up, Feed Back, and Feed Forward. It clearly identifies the goal of the passage, namely to motivate the topic convincingly, which is reflected in the score for Learning Goals (Score 2). It also explains what is missing in the current text by pointing out the unsupported claim and the need for a citation, corresponding to positive scores across the Feed Back criteria. Finally, it provides forward-looking guidance by recommending that important factual claims should be supported with evidence and statistics in future revisions, which is reflected in the scores for Self-regulation and Learning Skills (Score 1 each).
Example 1: Excerpt (Criteria: Anker) Draft Exposé "The modern museum has multiple purposes - to curate and preserve, to research, and to reach out to the public. They challenge us and ask us to question our assumptions about the past or the world around us." [1] […] But espacially the part about reaching out to the public doesn’t hold truth nowadays. Since the COVID-19 pandemic, museums worldwide have faced a dramatic decline in visitor numbers, highlighting the need for more engaging and innovative approaches to attract audiences. [Weakness: Where does this information come from? Cite.] Despite their [Weakness: Unclear reference; use a precise noun phrase such as “the museums”.] efforts to adapt many institutions still rely on traditional methods of interaction [Weakness: Underline why improving human interaction is important.], by simply showcasing an object and additional information written on text panels.[…] Review 1 (Junior Instructor) […]The anchor section provides a general context for the problem. However, the problem lacks detailed evidence or why it is important to research on it; for instance, while it mentions a "dramatic decline" since COVID-19, no quantitative data or case studies are included to support the claim. To strengthen the anchor, it is recommended to reference studies on the impact of the pandemic on cultural institutions and to underline its relevance/importance.[…] Review 2 (Student) […]In your motivation, you wrote that since the COVID-19 pandemic, museums faced a dramatic decline in visitor numbers. You aim to motivate your topic by providing an up-to-date example and an alarming development. This is important to make the reader clear why you think that this topic is relevant.[Feed Up] To ensure you provide valid and important facts, you should always provide a citation or source of such claims. This ensures that the reader will trust your judgment and understand the importance of your work. In this case, you do not provide any source or citation for the claim. For a sophisticated paper, you should avoid having unproofed claims in your work.[Feed Back] I would suggest citing some paper that proves your claim with statistics. This will lead to a better, more understandable paper and avoid having false claims in your work, which could be an undesired situation.[Feed Forward] […] Final Exposé […]But especially the part about reaching out to the public doesn’t hold truth nowadays. Since the COVID-19 pandemic, museums worldwide have faced a dramatic decline in visitor numbers [2], highlighting the need for more engaging and innovative approaches to attract audiences. For example, the National Gallery in London still had a 55% drop in visitors two years after the pandemic [3]. To counter this decline, museums need to innovate and create more engaging ways to interact with their content, because meaningful interaction is the key driver for museum visits. Despite the efforts to adapt, many institutions still rely on traditional methods of interaction[…]
Example 2: Teaser (Research Question)
This example shows how reviewers use the teaser criterion to assess whether a proposal states a clear research question. In the draft, the overall research direction is understandable, but the teaser remains too broad because it describes goals and anticipated contributions rather than a distinct research question. This is reflected especially clearly in the human reviews: one reviewer assigned a score of 1 and noted that a clearly formulated research question would give the exposé more structure, while another reviewer assigned a score of 0 and criticized the absence of a proper research question altogether. The automatic scoring of the draft points to the same central issue. First, several models took a strict view and penalized the text because no explicit research question was stated. For example, GPT 5.2 assigned a score of 0 and wrote that “No explicit research question is stated.” Second, some models interpreted the passage more leniently and treated it as an implicit research direction, while still criticizing the lack of precision. For instance, Claude Opus 4.6 assigned a score of 1 and argued that the work was described in broad terms without a clearly articulated research question. Third, a smaller group of models inferred a sufficiently clear question from the surrounding context. For example, Llama 3.3 assigned a score of 2 and described the research question as clearly formulated and aligned with the problem description.
Example 2: Excerpt (Criteria: Teaser (RQ)) Draft Exposé […]This research addresses the critical gap in current AI-driven cancer research by integrating the impact of the tumor microenvironment into data analysis workflows for personalized treatment planning. [Weakness: This states a research aim rather than an explicit research question.] We will discuss the landscape of adapting current AI-enhanced data analysis to the importance of TME, address the limitations of current AI models in incorporating TME data and explore the anticipated improvement of personalized treatment prediction. [Weakness: The goals remain broad; the main research question is not clearly stated.][…] Review 1 (Junior Instructor) Score: 1 […]Having an explicit, clearly defined research question gives a lot of structure to the overall expose, especially to the methodology section as your methods should tackle all details of all RQs. As you describe your work instead, the exact goals and your main focus are not entirely clear to me. When making your RQ more explicit, also keep in mind to check whether your methodology addresses all relevant sub-parts of your RQ(s). […] Review 2 (Senior Instructor) Score: 0 […]Also, a major issue is the lack of a distinct research question. The Motivation raises a question right at the end of the first paragraph. However, this question is rather a rhetorical question than a research question. Authors should include at least one sound research question that is answered in the work. […] Final Exposé Score: 2 […]The research will explore the following key questions: How does the heterogeneity of the tumor microenvironment impact treatment response predictions? What are the limitations of current AI-driven models in incorporating TME data, and how can these be addressed? How can integrating TME characteristics improve the predictive accuracy of personalized cancer treatments? […]
Appendix I Experimental Details
@@
We evaluate a diverse set of recent instruction-tuned LLMs spanning different architectures, scales, and deployment regimes. Our open-weight on-premise selection includes (i) a large dense decoder-only Transformer from the Llama 3.3 family (meta-llama/Llama-3.3-70B-Instruct; 70B parameters) Team (2024), (ii) a high-sparsity Mixture-of-Experts (MoE) model with long-context support (Qwen/Qwen3-Next-80B-A3B-Instruct; 80B total parameters with 3B active) Team (2025), as well as an independent MoE reasoning model in two variants, (iii) a high-capacity reasoning model (openai/gpt-oss-120b; 117B total parameters with 5.1B active) and (iv) a smaller low-latency variant (openai/gpt-oss-20b; 21B total parameters with 3.6B active) OpenAI (2025). These open-weight models can be executed locally on two NVIDIA A100 80GB GPUs, which is important for classroom deployment where data protection and limited consent often motivate on-premise inference rather than cloud-based APIs.
To contextualize results against strong proprietary baselines, we additionally include three closed-source, API-accessed models: OpenAI GPT 5.2 (openai/gpt-5.2), Anthropic Claude Opus 4.6 (anthropic/claude-opus-4-6), and Google Gemini 3.1 Pro Preview (gemini/gemini-3.1-pro-preview). Closed-source models are queried via the LiteLLM [^17] https://www.litellm.ai interface, enabling a unified prompting and logging pipeline across providers.
Inference for open-weight models is performed using the Hugging Face Transformers library [^18] https://huggingface.co/docs/transformers with deterministic greedy decoding (do_sample=False). For API-based models, we likewise use greedy decoding by default (temperature =0=0; and where supported, $\texttt{top\_p}=1$).
@@
For both tasks, we use a unified prompting strategy (Appendix J) with a system message that (i) assigns the model the role of an expert evaluator and (ii) enumerates the relevant criteria, including their names, short descriptions, allowed integer score ranges, and the verbal descriptions associated with each score level (Tables LABEL:tab:expose_rubric and LABEL:tab:review_rubric). The model is required to return a JSON list with one object per criterion, where each object contains the fields criterion_name, assigned_score, and justification. The additional justifications are requested to explain the decision-making process, as it is crucial in an environment where people are affected by the LLM outcome Kolářová and Schmude (2026).
@@
Appendix J Prompt Templates
@@
J.1 Expose Scoring Prompt
@@
J.2 Review Scoring Prompt
@@
Appendix K Additional Experimental Results
Criterion-wise alignment
Figures 7 and 8 report weighted mean binned JSD at the criterion level (darker = lower/better), complementing the domain-level averages in Table 25. Across both tasks, criteria on the 0–1–2 scale show a consistently higher JSD baseline than 0–1 criteria (Table 25), reflecting the increased degrees of freedom and typically higher annotator dispersion in three-way ratings, which makes distributional matching more challenging. For exposé, most criteria exhibit low divergence, but mismatches concentrate in the more interpretive rubric components—especially SOTA- and methodology-related criteria—with Anker emerging as uniformly difficult and Schedule Realistic Relevance showing a clear spike for some model/prompt settings. Reviews are more homogeneous overall: stylistic dimensions (language, tone) remain low across models, whereas divergence is driven primarily by epistemic “knowledge” judgments and reference-related criteria (notably Missing Reference).
Opus-4.6 Gem-3.1P GPT-5.2 OSS-120 OSS-20 L3.3-70 Qwen-80
Domain All Ind All Ind All Ind All Ind All Ind All Ind All Ind
Exposé criterion-based scoring (over 31 criteria)
0-1 (#21) 0.06 0.07 0.08 0.06 0.05 0.08 0.09 0.07 0.10 0.07 0.14 0.16 0.15 0.15
0-1-2 (#10) 0.08 0.09 0.22 0.23 0.10 0.13 0.09 0.21 0.10 0.26 0.24 0.35 0.24 0.26
All (#31) 0.07 0.08 0.13 0.12 0.07 0.10 0.09 0.11 0.10 0.13 0.17 0.22 0.18 0.19
Review criterion-based scoring (over 22 criteria)
0-1 (#10) 0.05 0.06 0.04 0.03 0.04 0.11 0.12 0.19 0.19 0.22 0.07 0.06 0.07 0.07
0-1-2 (#7) 0.07 0.12 0.12 0.10 0.24 0.19 0.09 0.10 0.08 0.09 0.11 0.11 0.21 0.21
-1-0 (#5) 0.03 0.10 0.03 0.03 0.03 0.14 0.07 0.11 0.06 0.11 0.03 0.16 0.04 0.08
All (#22) 0.05 0.09 0.06 0.06 0.10 0.14 0.10 0.14 0.13 0.15 0.08 0.10 0.11 0.12
Table 25: Weighted-mean Jensen–Shannon divergence (JSD) between human-vote and LLM-vote distributions, grouped by domain. Prompting variants: All, Ind. Lower is better. Model abbreviations: Opus-4.6=Claude Opus 4.6; Gem-3.1P=Gemini 3.1 Pro; GPT-5.2=GPT 5.2; OSS-120=GPT OSS (120B); OSS-20=GPT OSS (20B); L3.3-70=Llama 3.3 (70B); Qwen-80=Qwen 3 (80B).
Figure 6: Runtime per submission for criterion-based assessments across open-weight models. Bars show mean seconds per submission with standard deviation bars. Within each task, All denotes combined prompting (light fill) and Ind denotes single-criterion prompting (dark fill). Review runs are hatched; Exposé runs are unhatched. Ratios show $\times(\mathrm{Ind}/\mathrm{All})$, highlighting the additional latency introduced by single-criterion prompting.
@@
Across all models and both tasks, combined prompting is more efficient in most cases than single-criterion prompting: it reduces mean runtime per submission by about $1.3\times$$5.3\times$ and avoids the linear cost growth from issuing one model call per criterion (Figure 6). Single-criterion prompting also substantially increases total token processing because the full input is repeated across prompts, with particularly large overhead for long exposés and the review setting. These efficiency differences become especially consequential under the observed course usage dynamics. As shown by the behavioural interaction logs (Appendix Figure 5), student activity is strongly deadline-driven, with a pronounced spike in actions immediately before the review deadline, whereas instructor activity is relatively steady. This implies that in practice, LLM-based scoring must be provisioned for short periods of peak demand rather than average-day usage. As such, the overhead of single-criterion prompting is particularly problematic, while combined prompting better supports batching and throughput during deadline-induced bursts. Consequently, combined prompting is the recommended operating point for classroom deployment under constrained on-premise resources, with model choice primarily determining the remaining latency–quality trade-off. Runtime and token usage are reported in Appendix L (Tables 2627).